# Energy landscape analysis based on the Ising model: Tutorial review

### v1

@@ 2.5 Disconnectivity graph
The disconnectivity graph is a first-hand and main summary of the energy landscape. The disconnectivity graph is a tool from chemical physics and represents the relationship between different local minima of energy [59, 64]. For expository purposes, we show a hypothetical disconnectivity graph in Fig. 3(e); see Fig. 7(a) for the disconnectivity graph for a real fMRI data set from the human brain.
@@
Figure 7: Disconnectivity graphs. (a) Disconnectivity graph from fMRI data obtained from the brain of healthy human participants. We used the same data as the one used for Figure A1(a) in [32]. It is a concatenation of the first four fMRI recording sessions (resulting in $t_{\max}=818\times 4=3,272$) from a single participant in the Midnight Scan Club data set [65]. We used the first participant (i.e., MSC01). We used the so-called whole-brain network with $N=7$ nodes; the $N$ variables are the attention network (ATN), auditory network, default mode network (DMN), fronto-parietal network (FPN), salience network (SAN), somatosensory and motor network (SMN), and visual network [14]. See [32] for more details and the procedures for preprocessing the fMRI data. The activity pattern at each local minimum is indicated below the local minimum; a gray square represents $\sigma_{i}=-1$, and a white square represents $\sigma_{i}=1$. The accuracy of fit $r_{D}=0.915$. (b) Disconnectivity graph for randomly generated data with the same $N$ and $t_{\max}$ as those used in (a). (c) Disconnectivity graph obtained via pruning of (a). Local minima $\bm{\alpha}_{1}$ and $\bm{\alpha}_{5}$ in the disconnectivity graph shown in (a) have survived the significance test, whereas the other three local minima have not.
@@
There may be multiple local minima in the network of activity patterns. In the hypothetical network of activity patterns for $N=5$ variables shown in Fig. 8(a), we indicate the local minima by red circles. To construct the disconnectivity graph, we need to determine the energy barrier separating each pair of local minima. This task would be easy if the energy landscape were one-dimensional, as shown in Fig. 1(a) and (b) because there is only one shortest (i.e., non-redundant) path connecting the two local minima. In fact, this task is more involved because the actual energy landscape of the PMEM is defined on the hypercube. Let us consider possible paths between each pair of local minima. Partly because such a path does not have to be a shortest path (whose length is equal to the Hamming distance between the two local minima), there are many paths unless $N$ is small. The non-black lines in Fig. 8(a) show three among many paths connecting local minima $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$. Because the Hamming distance between $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$ is three, any path has at least length 3. For example, the path shown in green has length 5.
@@
Then, to determine the ease with which the system’s state can travel from $\bm{\alpha}_{1}$ to $\bm{\alpha}_{2}$ and vice versa, let us tabulate all the paths between them and the energy value of each node on the path. One can omit walks visiting the same node more than once because such redundant walks do not make the travels between $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$ easier than the corresponding shortest path. Figure 8(b) shows the three paths shown in color in Fig. 8(a) together with the energy on the nodes. The energy value of each node on the path dictates how much height of the hill the system’s state needs to climb to go from one end to the other. This is because, by definition, a high-energy activity pattern is a rare activity pattern in the data or the estimated PMEM. For example, in the path shown in skyblue, the highest hill to be overcome is activity pattern $\bm{\sigma}$, which has energy $E=-0.8$. The magenta path makes the travel easier with the largest energy of $E=-1.2$. Although the green path is longer than the skyblue path, the green path has the largest energy of $E=-1.1$, which is smaller than that for the skyblue path. As this example shows, longer paths may be easier to go through than shorter paths. If all the other paths between $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$ have an activity pattern with $E>-1.2$, we conclude that the magenta path is the easiest path between $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$ and record its peak height, denoted by $\overline{E}_{1,2}$ ($=-1.2$ in the current example). Computationally, we do not need to actually enumerate all the possible paths because there are efficient algorithms to realize the same computation [66, 6, 54, 32].
@@
How easy it is for the system’s state to travel between $\bm{\alpha}_{i}$ and $\bm{\alpha}_{j}$ depends on the energy of $\bm{\alpha}_{i}$ and $\bm{\alpha}_{j}$ as well as on $\overline{E}_{i,j}$. If $E(\bm{\alpha}_{i})$ is much smaller than $\overline{E}_{i,j}$, then it is difficult for a trajectory starting at $\bm{\alpha}_{i}$ to reach $\overline{E}_{i,j}$ even if $\overline{E}_{i,j}$ is small. We define the energy barrier between $\bm{\alpha}_{i}$ and $\bm{\alpha}_{j}$ from the viewpoint of $\bm{\alpha}_{i}$ by $\overline{E}_{i,j}-E(\bm{\alpha}_{i})$. The energy barrier between the same pair of local minima from the viewpoint of $\bm{\alpha}_{j}$, which is equal to $\overline{E}_{i,j}-E(\bm{\alpha}_{j})$, is different from that from the viewpoint of $\bm{\alpha}_{i}$ in general. For example, Fig. 8(b) indicates that $\overline{E}_{1,2}-E(\bm{\alpha}_{1})=-1.2-(-2.1)=0.9$ and $\overline{E}_{1,2}-E(\bm{\alpha}_{2})=-1.2-(-1.6)=0.4$.
@@
We carry out the same computation between each pair of local minima to obtain all $\overline{E}_{i,j}$. Once this is done, one can visualize all the local minima and energy barrier between each pair of them in a dendrogram, which is the disconnectivity graph, as shown in Figs. 3(e) and 7(a).
@@
Figure 8: Local minima and paths between them. (a) The network of activity patterns in the case of $N=5$ variables. The red circles represent local minima. (b) Three paths between local minima $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$ in the network of activity patterns shown in (a).
@@
The vertical axis of the disconnectivity graph represents the energy. A leaf, which vertically hangs down in the upside-down tree, represents a local minimum. The vertical position of the tip of the leaf represents the energy of the local minimum. Observe in Fig. 3(e) that $E(\bm{\alpha}_{1})=-2.1$ and $E(\bm{\alpha}_{2})=-1.6$, which are consistent with Fig. 8(b). We also observe that $\bm{\alpha}_{4}$ is the global minimum activity pattern in terms of the energy. Therefore, $\bm{\alpha}_{4}$ is the most frequent in the estimated PMEM and also likely so in the given data. The horizontal axis of the disconnectivity graph does not bear any meaning; different local minima are horizontally spread out only for visualization purposes. We also observe in Fig. 3(e) that the branch of $\bm{\alpha}_{1}$ and that of $\bm{\alpha}_{2}$ joins at $\overline{E}_{1,2}=-1.2$, which is also consistent with Fig. 8(b). In this manner, the disconnectivity graph represents all the local minima with their energy and the minimal height of the hill that any trajectory between each pair of local minima has to overcome.
@@
The disconnectivity graph shown in Fig. 3(e) suggests grouping of the six local minima into three groups, one composed of $\bm{\alpha}_{1}$, $\bm{\alpha}_{2}$, and $\bm{\alpha}_{3}$, one composed solely of $\bm{\alpha}_{4}$, and the other composed of $\bm{\alpha}_{5}$ and $\bm{\alpha}_{6}$. Alternatively, one may want to avoid grouping $\bm{\alpha}_{1}$ together with $\bm{\alpha}_{2}$ and $\bm{\alpha}_{3}$. One obtains such a grouping by cutting the disconnectivity graph into subtrees by drawing an arbitrary horizontal line into the disconnectivity graph (e.g., at $E=-1.1$, as shown by the dotted line in Fig. 3(e)). For example, the authors of [6] similarly grouped 10 local minima into three groups (shown in Fig. 1(e) of [6]). This procedure is the same as determination of the number of clusters in hierarchical clustering and necessarily arbitrary. Many studies relied on visual inspections or arbitrary criteria. We will explain a more principled method of grouping together some local minima into one major state in section 3.2.2.
@@ 2.6 Basin graph
The disconnectivity graph only contains the local minimum activity patterns. How should we interpret activity patterns that are not local minima? The low-dimensional schematics shown in Figs. 1 and 2 suggest that activity patterns that are not local minima belong to the basin of a local minimum. In other words, if we release a ball at the position of an arbitrary activity pattern, $\bm{\sigma}$, on the energy landscape, and if the ball only goes downhill (which is contrary to the actual stochastic dynamics implied by the PMEM though), the ball eventually reaches one of the local minimum $\bm{\alpha}$ and stops there. In this case, we say that $\bm{\sigma}$ belongs to the basin of $\bm{\alpha}$. In reality, visual understanding, such as in Figs. 1 and 2, is not easy to obtain because the network of activity patterns is a hypercube. However, the definition of the basin remains the same. We can tell the basin to which an arbitrary activity pattern, $\bm{\sigma}$, belongs as follows. We start by identifying the neighbor $\bm{\sigma}^{\prime}$ of $\bm{\sigma}$ that attains the lowest energy among all the $N$ neighbors of $\bm{\sigma}$ (see Fig. 9(a) for a schematic). Unless $\bm{\sigma}$ is itself a local minimum, $\bm{\sigma}^{\prime}$ has a lower energy than $\bm{\sigma}$, which we assume here. If there is a tie, we break it arbitrarily to set $\bm{\sigma}^{\prime}$. The direction from $\bm{\sigma}$ to $\bm{\sigma}^{\prime}$ is the steepest descent direction from $\bm{\sigma}$. Therefore, we move from $\bm{\sigma}$ and $\bm{\sigma}^{\prime}$ in search of the bottom of the basin. Then, at $\bm{\sigma}^{\prime}$, we similarly identify the neighboring activity pattern having the lowest energy, denoted by $\bm{\sigma}^{\prime\prime}$. If $\bm{\sigma}^{\prime\prime}$’s energy is lower than $\bm{\sigma}^{\prime}$’s energy, then we move to $\bm{\sigma}^{\prime\prime}$. Otherwise, $\bm{\sigma}^{\prime}$ has an energy that is lower than all its neighbors such that $\bm{\sigma}^{\prime}$ is a local minimum. In the latter case, we stop the journey at $\bm{\sigma}^{\prime}$ and conclude that $\bm{\sigma}$ belongs to the basin of $\bm{\sigma}^{\prime}$. If this is not the case, we repeat the procedure until we reach a local minimum, and we call that $\bm{\sigma}$ belongs to the basin of the finally visited activity pattern, which is a local minimum. In Fig. 9(a), the steepest descent path starting at $\bm{\sigma}$ stops at $\bm{\sigma}^{\prime\prime}$. Therefore, $\bm{\sigma}^{\prime\prime}$ is the local minimum associated with $\bm{\sigma}$. In this manner, we can identify to which basin each activity pattern belongs.
@@
Figure 9: Basin graph. (a) The steepest descent path from $\bm{\sigma}$ to its associated local minimum $\bm{\sigma}^{\prime\prime}$. A node represents an activity pattern. The numbers attached to each activity pattern represents its energy value. We set $N=5$. In fact, $\bm{\sigma}$ and $\bm{\sigma}^{\prime\prime}$ share two neighbors because they are distance 2 apart from each other. One such neighbor is $\bm{\sigma}^{\prime}$. The other shared neighbor is an activity pattern with $E=-0.6$. For expository purposes, we showed this second neighbor as separate nodes in this figure. (b) An empirical basin graph with $N=7$ corresponding to the disconnectivity graph shown in Fig. 7(a).
@@
What we call the basin graph tells us the basin of each activity pattern and more; see Fig. 3(f) for a schematic and Fig. 9(b) for a real example. A node of the basin graph is an activity pattern. Therefore, there are $2^{N}$ nodes; in Figs. 3(f) and 9(b), there are $322^{5}=$ 32 and $1282^{7}=$ 128 nodes, respectively. The edge of the basin graph represents the steepest descent move from an activity pattern to its neighbor. Differently from the network of activity patterns, which is also a network among the $2^{N}$ activity pattern nodes, the edge of the basin graph is directed, and each node has at most one out-going edge. If a node is not a local minimum, it has one out-going edge (i.e., the node’s out-degree is equal to 1); see the orange nodes in Figs. 3(f) and 9(b). If a node is a local minimum, it does not have any out-going edge; see the blue nodes in the same figures. Any node has between 0 to $N$ in-coming edges; its number is called the in-degree of the node. The basin graph is composed of connected components (i.e., islands), each of which contains just one local minimum. This fact aids the interpretation of the basin graph. The number of nodes in the connected components is equal to the size of the basin of the corresponding local minimum. Comparison between Figs. 3(e) and 3(f) and between Figs. 7(a) and 9(b) suggests that an local minimum with a smaller energy value (i..e, therefore, high frequency of appearance) tends to have a large basin. However, while this association is often found in empirical data, it is not guaranteed.
@@
Some activity patterns may be located near the border between different basins. Such an activity pattern roughly, but not exactly, corresponds to those on the top of a hill separating two basins (e.g., the local maximum in Fig. 1 or a dotted line in Fig. 2). Precisely speaking, if activity pattern $\bm{\sigma}$ belonging to the basin of $\bm{\alpha}_{i}$ has a neighboring activity pattern $\bm{\sigma}^{\prime}$ belonging to the basin of a different local minimum, then we say that $\bm{\sigma}$ and $\bm{\sigma}^{\prime}$ are at the boundary between the two basins.

### v2

@@ 2.5 Disconnectivity graph
The disconnectivity graph is a first-hand and main summary of the energy landscape. The disconnectivity graph is a tool from chemical physics and represents the relationship between different local minima of energy [59, 65]. For expository purposes, we show a hypothetical disconnectivity graph in Fig 3(e); see Fig 7(a) for the disconnectivity graph for a real fMRI data set from the human brain.
@@
Fig 7: Disconnectivity graphs. (a) Disconnectivity graph from fMRI data obtained from the brain of healthy human participants. We used the same data as the one used for Fig A1(a) in [32]. It is a concatenation of the first four fMRI recording sessions (resulting in $t_{\max}=818\times 4=3,272$) from a single participant in the Midnight Scan Club data set [64]. We used the first participant (i.e., MSC01). We used the so-called whole-brain network with $N=7$ nodes; the $N$ variables are the attention network (ATN), auditory network, default mode network (DMN), fronto-parietal network (FPN), salience network (SAN), somatosensory and motor network (SMN), and visual network [14]. See [32] for more details and the procedures for preprocessing the fMRI data. The activity pattern at each local minimum is indicated below the local minimum; a gray square represents $\sigma_{i}=-1$, and a white square represents $\sigma_{i}=1$. The accuracy of fit $r_{D}=0.915$. (b) Disconnectivity graph for randomly generated data with the same $N$ and $t_{\max}$ as those used in (a). (c) Disconnectivity graph obtained via pruning of (a). Local minima $\bm{\alpha}_{1}$ and $\bm{\alpha}_{5}$ in the disconnectivity graph shown in (a) have survived the significance test, whereas the other three local minima have not.
@@
There may be multiple local minima in the network of activity patterns. In the hypothetical network of activity patterns for $N=5$ variables shown in Fig 8(a), we indicate the local minima by red circles. To construct the disconnectivity graph, we need to determine the energy barrier separating each pair of local minima. This task would be easy if the energy landscape were one-dimensional, as shown in Fig 1(a) and (b) because there is only one shortest (i.e., non-redundant) path connecting the two local minima. In fact, this task is more involved because the actual energy landscape of the PMEM is defined on the hypercube. Let us consider possible paths between each pair of local minima. Partly because such a path does not have to be a shortest path (whose length is equal to the Hamming distance between the two local minima), there are many paths unless $N$ is small. The non-black lines in Fig 8(a) show three among many paths connecting local minima $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$. Because the Hamming distance between $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$ is three, any path has at least length 3. For example, the path shown in green has length 5.
@@
Then, to determine the ease with which the system’s state can travel from $\bm{\alpha}_{1}$ to $\bm{\alpha}_{2}$ and vice versa, let us tabulate all the paths between them and the energy value of each node on the path. One can omit walks visiting the same node more than once because such redundant walks do not make the travels between $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$ easier than the corresponding shortest path. Fig 8(b) shows the three paths shown in color in Fig 8(a) together with the energy on the nodes. The energy value of each node on the path dictates how much height of the hill the system’s state needs to climb to go from one end to the other. This is because, by definition, a high-energy activity pattern is a rare activity pattern in the data or the estimated PMEM. For example, in the path shown in skyblue, the highest hill to be overcome is activity pattern $\bm{\sigma}$, which has energy $E=-0.8$. The magenta path makes the travel easier with the largest energy of $E=-1.2$. Although the green path is longer than the skyblue path, the green path has the largest energy of $E=-1.1$, which is smaller than that for the skyblue path. As this example shows, longer paths may be easier to go through than shorter paths. If all the other paths between $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$ have an activity pattern with $E>-1.2$, we conclude that the magenta path is the easiest path between $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$ and record its peak height, denoted by $\overline{E}_{1,2}$ ($=-1.2$ in the current example). Computationally, we do not need to actually enumerate all the possible paths because there are efficient algorithms to realize the same computation [66, 6, 54, 32].
@@
How easy it is for the system’s state to travel between $\bm{\alpha}_{i}$ and $\bm{\alpha}_{j}$ depends on the energy of $\bm{\alpha}_{i}$ and $\bm{\alpha}_{j}$ as well as on $\overline{E}_{i,j}$. If $E(\bm{\alpha}_{i})$ is much smaller than $\overline{E}_{i,j}$, then it is difficult for a trajectory starting at $\bm{\alpha}_{i}$ to reach $\overline{E}_{i,j}$ even if $\overline{E}_{i,j}$ is small. We define the energy barrier between $\bm{\alpha}_{i}$ and $\bm{\alpha}_{j}$ from the viewpoint of $\bm{\alpha}_{i}$ by $\overline{E}_{i,j}-E(\bm{\alpha}_{i})$. The energy barrier between the same pair of local minima from the viewpoint of $\bm{\alpha}_{j}$, which is equal to $\overline{E}_{i,j}-E(\bm{\alpha}_{j})$, is different from that from the viewpoint of $\bm{\alpha}_{i}$ in general. For example, Fig 8(b) indicates that $\overline{E}_{1,2}-E(\bm{\alpha}_{1})=-1.2-(-2.1)=0.9$ and $\overline{E}_{1,2}-E(\bm{\alpha}_{2})=-1.2-(-1.6)=0.4$.
@@
We carry out the same computation between each pair of local minima to obtain all $\overline{E}_{i,j}$. Once this is done, one can visualize all the local minima and energy barrier between each pair of them in a dendrogram, which is the disconnectivity graph, as shown in Figs 3(e) and 7(a).
@@
Fig 8: Local minima and paths between them. (a) The network of activity patterns in the case of $N=5$ variables. The red circles represent local minima. (b) Three paths between local minima $\bm{\alpha}_{1}$ and $\bm{\alpha}_{2}$ in the network of activity patterns shown in (a).
@@
The vertical axis of the disconnectivity graph represents the energy. A leaf, which vertically hangs down in the upside-down tree, represents a local minimum. The vertical position of the tip of the leaf represents the energy of the local minimum. Observe in Fig 3(e) that $E(\bm{\alpha}_{1})=-2.1$ and $E(\bm{\alpha}_{2})=-1.6$, which are consistent with Fig 8(b). We also observe that $\bm{\alpha}_{4}$ is the global minimum activity pattern in terms of the energy. Therefore, $\bm{\alpha}_{4}$ is the most frequent in the estimated PMEM and also likely so in the given data. The horizontal axis of the disconnectivity graph does not bear any meaning; different local minima are horizontally spread out only for visualization purposes. We also observe in Fig 3(e) that the branch of $\bm{\alpha}_{1}$ and that of $\bm{\alpha}_{2}$ joins at $\overline{E}_{1,2}=-1.2$, which is also consistent with Fig 8(b). In this manner, the disconnectivity graph represents all the local minima with their energy and the minimal height of the hill that any trajectory between each pair of local minima has to overcome.
@@
The disconnectivity graph shown in Fig 3(e) suggests grouping of the six local minima into three groups, one composed of $\bm{\alpha}_{1}$, $\bm{\alpha}_{2}$, and $\bm{\alpha}_{3}$, one composed solely of $\bm{\alpha}_{4}$, and the other composed of $\bm{\alpha}_{5}$ and $\bm{\alpha}_{6}$. Alternatively, one may want to avoid grouping $\bm{\alpha}_{1}$ together with $\bm{\alpha}_{2}$ and $\bm{\alpha}_{3}$. One obtains such a grouping by cutting the disconnectivity graph into subtrees by drawing an arbitrary horizontal line into the disconnectivity graph (e.g., at $E=-1.1$, as shown by the dotted line in Fig 3(e)). For example, the authors of [6] similarly grouped 10 local minima into three groups (shown in Fig 1(e) of [6]). This procedure is the same as determination of the number of clusters in hierarchical clustering and necessarily arbitrary. Many studies relied on visual inspections or arbitrary criteria. We will explain a more principled method of grouping together some local minima into one major state in Section 3.2.2.
@@ 2.6 Basin graph
The disconnectivity graph only contains the local minimum activity patterns. How should we interpret activity patterns that are not local minima? The low-dimensional schematics shown in Figs 1 and 2 suggest that activity patterns that are not local minima belong to the basin of a local minimum. In other words, if we release a ball at the position of an arbitrary activity pattern, $\bm{\sigma}$, on the energy landscape, and if the ball only goes downhill (which is contrary to the actual stochastic dynamics implied by the PMEM though), the ball eventually reaches one of the local minimum $\bm{\alpha}$ and stops there. In this case, we say that $\bm{\sigma}$ belongs to the basin of $\bm{\alpha}$. In reality, visual understanding, such as in Figs 1 and 2, is not easy to obtain because the network of activity patterns is a hypercube. However, the definition of the basin remains the same. We can tell the basin to which an arbitrary activity pattern, $\bm{\sigma}$, belongs as follows. We start by identifying the neighbor $\bm{\sigma}^{\prime}$ of $\bm{\sigma}$ that attains the lowest energy among all the $N$ neighbors of $\bm{\sigma}$ (see Fig 9(a) for a schematic). Unless $\bm{\sigma}$ is itself a local minimum, $\bm{\sigma}^{\prime}$ has a lower energy than $\bm{\sigma}$, which we assume here. If there is a tie, we break it arbitrarily to set $\bm{\sigma}^{\prime}$. The direction from $\bm{\sigma}$ to $\bm{\sigma}^{\prime}$ is the steepest descent direction from $\bm{\sigma}$. Therefore, we move from $\bm{\sigma}$ and $\bm{\sigma}^{\prime}$ in search of the bottom of the basin. Then, at $\bm{\sigma}^{\prime}$, we similarly identify the neighboring activity pattern having the lowest energy, denoted by $\bm{\sigma}^{\prime\prime}$. If $\bm{\sigma}^{\prime\prime}$’s energy is lower than $\bm{\sigma}^{\prime}$’s energy, then we move to $\bm{\sigma}^{\prime\prime}$. Otherwise, $\bm{\sigma}^{\prime}$ has an energy that is lower than all its neighbors such that $\bm{\sigma}^{\prime}$ is a local minimum. In the latter case, we stop the journey at $\bm{\sigma}^{\prime}$ and conclude that $\bm{\sigma}$ belongs to the basin of $\bm{\sigma}^{\prime}$. If this is not the case, we repeat the procedure until we reach a local minimum, and we call that $\bm{\sigma}$ belongs to the basin of the finally visited activity pattern, which is a local minimum. In Fig 9(a), the steepest descent path starting at $\bm{\sigma}$ stops at $\bm{\sigma}^{\prime\prime}$. Therefore, $\bm{\sigma}^{\prime\prime}$ is the local minimum associated with $\bm{\sigma}$. In this manner, we can identify to which basin each activity pattern belongs.
@@
Fig 9: Basin graph. (a) The steepest descent path from $\bm{\sigma}$ to its associated local minimum $\bm{\sigma}^{\prime\prime}$. A node represents an activity pattern. The numbers attached to each activity pattern represents its energy value. We set $N=5$. In fact, $\bm{\sigma}$ and $\bm{\sigma}^{\prime\prime}$ share two neighbors because they are distance 2 apart from each other. One such neighbor is $\bm{\sigma}^{\prime}$. The other shared neighbor is an activity pattern with $E=-0.6$. For expository purposes, we showed this second neighbor as separate nodes in this figure. (b) An empirical basin graph with $N=7$ corresponding to the disconnectivity graph shown in Fig 7(a).
@@
What we call the basin graph tells us the basin of each activity pattern and more; see Fig 3(f) for a schematic and Fig 9(b) for a real example. A node of the basin graph is an activity pattern. Therefore, there are $2^{N}$ nodes; in Figs 3(f) and 9(b), there are $322^{5}=$ 32 and $1282^{7}=$ 128 nodes, respectively. The edge of the basin graph represents the steepest descent move from an activity pattern to its neighbor. Differently from the network of activity patterns, which is also a network among the $2^{N}$ activity pattern nodes, the edge of the basin graph is directed, and each node has at most one out-going edge. If a node is not a local minimum, it has one out-going edge (i.e., the node’s out-degree is equal to 1); see the orange nodes in Figs 3(f) and 9(b). If a node is a local minimum, it does not have any out-going edge; see the blue nodes in the same figures. Any node has between 0 to $N$ in-coming edges; its number is called the in-degree of the node. The basin graph is composed of so-called weakly connected components (i.e., islands), each of which contains just one local minimum. This fact aids the interpretation of the basin graph. The number of nodes in the connected components is equal to the size of the basin of the corresponding local minimum. Comparison between Figs 3(e) and 3(f) and between Figs 7(a) and 9(b) suggests that an local minimum with a smaller energy value (i..e, therefore, high frequency of appearance) tends to have a large basin. However, while this association is often found in empirical data, it is not guaranteed.
@@
Some activity patterns may be located near the border between different basins. Such an activity pattern roughly, but not exactly, corresponds to those on the top of a hill separating two basins (e.g., the local maximum in Fig 1 or a dotted line in Fig 2). Precisely speaking, if activity pattern $\bm{\sigma}$ belonging to the basin of $\bm{\alpha}_{i}$ has a neighboring activity pattern $\bm{\sigma}^{\prime}$ belonging to the basin of a different local minimum, then we say that $\bm{\sigma}$ and $\bm{\sigma}^{\prime}$ are at the boundary between the two basins.